# Scaling Motion Forecasting Models with Ensemble Distillation

### v1

@@ IV-B Distillation losses
Minimizing this loss by varying $\boldsymbol{\Phi_{Student}}$ will maximize the student probability of the set of teacher samples. By doing so, the student model outputs become similar to that of the teacher. An additional loss term is added to increase the spikiness of the distribution by minimizing the entropy of the student mixture weights.
$\mathcal{L}_{DistillEntropy}=-\sum_{n=1}^{N}\pi^{n}log(\pi^{n})$ (7)
@@
$\mathcal{L}_{Total}=$ $\mathcal{L}_{DistillNLL}+w_{ent}\mathcal{L}_{DistillEntropy}+w_{gt}\mathcal{L}_{gt}$ (8)
where $w_{ent}$ is a weight for the entropy loss term and $\mathcal{L}_{gt}$ is the groundtruth loss used when training the model without distillation. We follow the same formulation as [33] for the groundtruth loss.
@@ V-A Datasets & Metrics
For experiments we use two datasets: (1) the Waymo Open Motion dataset (WOMD) [12] which contains over 570 hours of driving data including maps, traffic light states, and agent motion data consisting of 3D bounding boxes sampled at 10 Hz; (2) the Argoverse 2 Motion Forecasting dataset [42] which consists of 250,000 scenarios mined for challenging interactions between the autonomous vehicle and other actors in each local scene, including maps and agent motion data sampled at 10Hz. The WOMD task is to predict an 8 second future trajectory for up to 8 agents given a 1 second history of agent trajectories, whereas the Argoverse task is to predict a 6 second future trajectory for a single ‘focal’ agent per scenario given 5 seconds of past history.
@@ V-B Training Details and Hyperparameters
Models minFDE ($\downarrow$) minADE ($\downarrow$) MR ($\downarrow$) Overlap ($\downarrow$) mAP ($\uparrow$) $*{}^{*}$ ($\uparrow$) rel-FLOPs($\downarrow$)
@@ VI-A Ensembles
It performs particularly well on minADE and minFDE metrics relative to the state of the art. Our ensemble of $K=28$ teacher models achieves competitive performance for all metrics, placing second on the WOMD leaderboard only behind another large ensemble. Experiments on the Argoverse dataset shown in Fig 5 show a similar positive log-linear scaling relationship with FLOPs for ensembles where the orange dots and lines represent the results for the ensemble models. The 10 model ensemble places 3rd on the Argoverse leaderboard. These results importantly, validate the hypothesis that ensembling can provide a way to scale model performance on motion forecasting problems. We also analyze qualitative improvements to model outputs in Fig 6, with examples of prediction quality of the ensemble model.
@@ VI-B Ensemble Distillation
The green dots and lines in figure 4 show the WOMD validation metrics and trends of results for distillation models as a function of FLOPs. All distillation models are distilled from the $K=28$ model ensemble as described above using the efficient sampling procedure described in IV-D and setting $\tau=8$, $w_{gt}=0.4$, $w_{ent}=1.0$, and $w_{var}=0.1$. The first result to note is that the smallest distilled model (the far left green point) outperforms its corresponding baseline model with the same architecture (the far left orange point) by a large margin on all metrics while requiring far fewer FLOPs. This is due to the additional evidence per training example that we get when training models with distillation losses. While increasing parameters in the teacher models resulted in over-fitting, it is notable that we can increase parameters/FLOPs in the student model to improve performance. This due to the extra information provided by the teacher predictions. All metrics improve with increased size of the 2X scaled distilled student model. The distillation result for the Argoverse dataset is shown by the green dots and lines in fig 5. Here we see that the distilled models also perform far better than the base model. For the Argoverse experiments we trained one model with 0.75X of the FLOPs required by the base model and one with required FLOPs equal to the base model. Both models show significant performance improvement over the baseline model (the far left orange dot). Despite having 25% fewer required FLOPs, the smaller model still significantly outperforms the baseline model. The key result is that training a distilled model from a much larger ensemble provides a significant increase in performance with much lower compute cost than the ensemble, remaining suitable for running onboard the robot.
@@
Table I shows a comparison of our models with the current WOMD leaderboard. Our ensemble model ranks second on the leaderboard. The distilled model has reduced performance but with over 20X fewer FLOPs required. It is not known at this time how the other entries on the leaderboard compare to the distilled model in terms of required FLOPs. The key result is that the distilled model provides high performance in a small model that can be deployed in real-time onboard robotics applications.

### v2

@@ IV-B Distillation losses
Minimizing this loss by varying $\boldsymbol{\Phi_{Student}}$ will maximize the student probability of the set of teacher samples. By doing so, the student model outputs become similar to that of the teacher.
@@
$\mathcal{L}_{Total}=$ $\mathcal{L}_{DistillNLL}+w_{gt}\mathcal{L}_{gt}$ (7)
where $\mathcal{L}_{gt}$ is the groundtruth loss used when training the model without distillation. We follow the same formulation as [33] for the groundtruth loss.
@@ V-A Datasets & Metrics
For experiments we use two datasets: (1) the Waymo Open Motion dataset (WOMD) [12] which contains over 570 hours of driving data including maps, traffic light states, and agent motion data consisting of 3D bounding boxes sampled at 10 Hz; (2) the Argoverse 2 Motion Forecasting dataset [42] which consists of 250,000 scenarios mined for challenging interactions between the autonomous vehicle and other actors in each local scene, including maps and agent motion data sampled at 10Hz. The WOMD task is to predict an 8 second future trajectory for up to 8 agents given a 1 second history of agent trajectories, whereas the Argoverse task is to predict a 6 second future trajectory for a single ‘focal’ agent per scenario given 5 seconds of past history. For WOMD training, examples with classified u-turn and right and left turn trajectories were duplicated at a rate of 5% to improve performance on rarer trajectories.
@@ V-B Training Details and Hyperparameters
Models minFDE ($\downarrow$) minADE ($\downarrow$) MR ($\downarrow$) Overlap ($\downarrow$) mAP ($\uparrow$) soft-mAP ($\uparrow$) rel-FLOPs($\downarrow$)
@@ VI-A Ensembles
It performs particularly well on minADE and minFDE metrics relative to the state of the art. Our ensemble of $K=20$ teacher models achieves competitive performance for all metrics, placing third on the WOMD leaderboard. Future experiments will update this figure with distilled models with larger capacity. Experiments on the Argoverse dataset shown in Fig 5 show a similar positive log-linear scaling relationship with FLOPs for ensembles where the orange dots and lines represent the results for the ensemble models. The 10 model ensemble places 3rd on the Argoverse leaderboard. These results importantly, validate the hypothesis that ensembling can provide a way to scale model performance on motion forecasting problems. We also analyze qualitative improvements to model outputs in Fig 6, with examples of prediction quality of the ensemble model.
@@ VI-B Ensemble Distillation
The green dots and lines in figure 4 show the WOMD validation metrics and trends of results for distillation models as a function of FLOPs. All distillation models are distilled from the $K=20$ model ensemble as described above using the efficient sampling procedure described in IV-D and setting $\tau=8$, $w_{gt}=0.4$, and $w_{var}=0.5$. The first result to note is that the distilled model (the far left green point) outperforms its corresponding baseline model with the same architecture (the far left orange point) by a large margin on all metrics while requiring far fewer FLOPs. This is due to the additional evidence per training example that we get when training models with distillation losses. Notably, the minADE and miss rate metrics of the distilled model are better than the teacher ensemble. The distillation result for the Argoverse dataset is shown by the green dots and lines in fig 5. Here we see that the distilled models also perform far better than the base model. For the Argoverse experiments we trained one model with 0.75X of the FLOPs required by the base model and one with required FLOPs equal to the base model. Both models show significant performance improvement over the baseline model (the far left orange dot). Despite having 25% fewer required FLOPs, the smaller model still significantly outperforms the baseline model. The key result is that training a distilled model from a much larger ensemble provides a significant increase in performance with much lower compute cost than the ensemble, remaining suitable for running onboard the robot.
@@
Table I shows a comparison of our models with the current WOMD leaderboard. Our ensemble model ranks third on the leaderboard. The distilled model has reduced performance on most metrics but with over 20X fewer FLOPs required. It is not known at this time how the other entries on the leaderboard compare to the distilled model in terms of required FLOPs. Notably the distilled model has the best minADE performance of any model on the leaderboard. The key result is that the distilled model provides high performance in a small model that can be deployed in real-time onboard robotics applications.